Circular No. 7258

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE P. Astier, Laboratoire de Physique Nucléaire et des Hautes Energies,

CNRS/IN2P3, and Universités Paris VI et VII, on behalf of the European Supernova Cosmology Consortium (cf. IAUC 7182, plus A. Goobar, University of Stockholm), and in collaboration with the Supernova Cosmology Project (cf. IAUC 7128), report ten supernovae found with the 2.5-m Isaac Newton Telescope (INT) in their Intermediate Redshift Supernova Search:

SN  1999 UT  α_{2000}  δ_{2000}  g′  z Type
1999dr Sept. 2.99 23^{h}00^{m}17.^{s}56 − 0^{o}05′12′′.5 22.1 0.182 Ia
1999ds Sept. 7.06 23 53 51.63 + 0 09 22.9 22.4 0.350 II?
1999dt Sept. 4.05 0 45 42.29 + 0 03 22.2 23.5 0.439 Ia
1999du Sept. 8.16 1 07 05.94 − 0 07 53.8 22.8 0.263 Ia
1999dv Sept. 9.10 1 08 58.96 + 0 00 24.8 21.8 0.187 Ia
1999dw Sept. 7.15 1 22 52.80 − 0 16 20.8 24.1 0.447 Ia?
1999dx Sept. 8.20 1 33 59.45 + 0 04 15.3 22.2 0.270 Ia
1999dy Sept. 8.21 1 35 49.53 + 0 08 38.3 21.7 0.206 Ia
1999dz Sept. 8.21 1 37 03.24 + 0 01 57.9 23.4 0.488 Ia
1999ea Sept. 9.20 1 47 26.09 − 0 02 07.2 23.3 0.399 Ia

Most of the supernovae were discovered near maximum light and reobserved 1–2 days after discovery on the INT. Confirmation spectroscopy was obtained on the 4.2-m William Herschel Telescope during the nights of Sept. 12–14. The supernovae are all within 1′′.5 of the host galaxies’ centers. Further photometry in g′ and r′ is being obtained with various telescopes. Finding charts are available from hardin@snova1.in2p3.fr.

V1493 AQUILAE D. K. Lynch, R. J. Rudy, and S. Mazuk, The Aerospace Corporation;

and R. C. Puetter, Center for Astrophysics and Space Science, University of California at San Diego, report 0.8–2.5-μm spectrophotometry of V1493 Aql = N Aql 1999 using the Shane 3-m telescope at Lick Observatory (+ Aerospace Near Infrared Imaging Spectrograph) on Aug. 28.27 UT (∼ 48 days after outburst). The spectrum shows very broad lines (FWHM ∼ 4000 km/s), with the strongest being the Lyβ fluoresced O I lines at 0.8446 and 1.1287 μm. Also present were the O I lines at 1.3164 μm, excited by continuum fluorescence. The Paschen and Brackett lines were also strong, as was the He I at 1.0831 μm. There was no evidence of thermal emission from dust. Infrared magnitudes: J = 8.4, H = 8.2, K = 7.7.

1999 September 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green